The kinetic energy of an object of mass, m, moving with a speed “v” is 25J. By what factor will its kinetic energy change if its mass is tripled while the velocity is decreased by one third?
1
Expert's answer
2020-09-02T12:14:24-0400

Kinetic energy of an object is described by a formula "T=\frac{mv^2}{2}." Initial energy of the given object with mass "m_1" and velocity "v_1": "T_1=\frac{m_1v_1^2}{2}=25J." After we changed the mass and the velocity, we have:"m_2=3m_1, v_2=\frac{v_1}{3}". Thus, "T_2=\frac{m_2v_2^2}{2}=\frac{3m_1}{2}(\frac{v_1}{3})^2=\frac{m_1v_1^2}{6}=\frac{T_1}{3}." Hence, the kinetic energy will decrase by "\frac{1}{3}."